# captionUnknown document class (or package)

### v1

@@ Abstract
As large language models (LLMs) evolve, their integration with 3D spatial data (3D-LLMs) has seen rapid progress, offering unprecedented capabilities for understanding and interacting with physical spaces. This survey provides a comprehensive overview of the methodologies enabling LLMs to process, understand, and generate 3D data. Highlighting the unique advantages of LLMs, such as in-context learning, step-by-step reasoning, open-vocabulary capabilities, and extensive world knowledge, we underscore their potential to significantly advance spatial comprehension and interaction within embodied Artificial Intelligence (AI) systems. Our investigation spans various 3D data representations, from point clouds to Neural Radiance Fields (NeRFs). It examines their integration with LLMs for tasks such as 3D scene understanding, captioning, question-answering, and dialogue, as well as LLM-based agents for spatial reasoning, planning, and navigation. The paper also includes a brief review of other methods that integrate 3D and language. The meta-analysis presented in this paper reveals significant progress yet underscores the necessity for novel approaches to harness the full potential of 3D-LLMs. Hence, with this paper, we aim to chart a course for future research that explores and expands the capabilities of 3D-LLMs in understanding and interacting with the complex 3D world. To support this survey, we have established a project page where papers related to our topic are organized and listed: https://github.com/ActiveVisionLab/Awesome-LLM-3D.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has marked a transformative era in natural language processing, enabling machines to understand, generate, and interact with human language in previously unimagined ways. However, the physical world around us is inherently three-dimensional, and understanding spatial 3D environments is crucial for many real-world applications that involve perception, navigation, and interaction within these 3D spaces. With recent advancements, the application of LLMs has extended well beyond text. The fusion of LLMs with 3D data presents a unique opportunity to enhance computational models’ understanding of and interaction with the physical world, leading to innovations across various domains, including autonomous systems [1, 2, 3, 4, 5], augmented reality [6, 7, 8, 9], robotic navigation [10, 11, 12], and robotic manipulation [13, 14, 15].
Recent works have demonstrated the potential of integrating LLMs with 3D data to interpret, reason, or plan in complex 3D environments, by leveraging the inherent strengths of LLMs, including zero-shot learning [16, 17], advanced reasoning [13, 18, 19], and extensive knowledge [20, 21]. However, the integration of LLMs with 3D data is not trivial. Issues such as 3D data representation, model scalability, and computational efficiency remain significant hurdles. Moreover, ensuring the models can operate in real-world settings requires overcoming obstacles related to data diversity and environmental complexity. Addressing these challenges is crucial for the full realization of LLMs’ potential in 3D applications to create dynamic and context-aware AI systems.
This survey paper provides a critical examination of the intersection of LLMs with 3D data, offering an exhaustive overview of current methodologies, applications, and challenges in this domain. We start by providing the relevant background on common 3D representations, a brief introduction to LLMs and an overview of Vision-Language Models (VLMs) and Vision Foundation Models (VFMs) in Sec. 2. In Sec. 3, we detail the 3D vision language tasks that current methods aim to solve, outlining the current evaluation metrics and protocols.
Figure 1: A general timeline of methods leading up to Large Language Models (LLMs) in 3D.
Next, in Sec. 4, we analyze the data formats, processing techniques, and model architectures that have shown promise in enhancing 3D understanding through the lens of LLM capabilities. We showcase a variety of areas where the amalgamation of LLMs and 3D data has been successfully demonstrated, such as: using LLMs’ world-knowledge [20, 22] and reasoning capabilities [21, 23] to enhance 3D task performance, using LLMs as multi-modal interfaces [24, 18] and embodied agents [15, 13], or generating complex scenes with LLMs [25, 26]. Beyond LLMs, a few research works have proposed end-to-end architectures that unify 3D perception with language capabilities [27, 28]. Moreover, a wide range of works have explored distilling the knowledge of off-the-shelf 2D Vision Language Models (VLMs) for open vocabulary 3D scene understanding [29, 30] as well as text-driven 3D generation [31, 32]. This survey provides a complete overview of these methods in Sec. 5 to show a complete picture of the field of 3D+Language. We then outline the datasets used to train and evaluate these methods in Sec. 6. Finally, Sec. 7 highlights the challenges and potential future research directions in this field.
@@ 2 Background
This section offers essential background knowledge on 3D representations, Large Language Models (LLMs), 2D Vision-Language Models (VLMs), and Vision Foundation Models (VFMs).
@@
Choosing a 3D representation to describe, model, and understand our world is a crucial topic that contributes to the understanding of the current progress in 3D-LLMs. It is also a fundamental research area in computer vision. The field has recently seen substantial growth due to advancements in deep learning, computational resources, and 3D data availability. We briefly describe the most common 3D representations currently in use.
Point Clouds represent 3D shapes with a set of data points in space, storing each point position in a 3D Cartesian coordinate system. In addition to storing the position, other information (e.g. color, normal) can also be stored with each point. Point cloud-based methods [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] are noted for their low storage footprint but lack surface topology information. Typical sources for obtaining point clouds include LiDAR sensors, structured light scanners, time-of-flight cameras, stereo-views, photogrammetry, etc.
Voxel Grids [44, 45, 46, 47, 48, 49, 50, 51] consist of unit cubes in 3D space, akin to pixel representations in 2D [52]. Each voxel minimally encodes the occupancy information (binary or probabilistically) but can additionally encode distance to a surface, as in the Signed Distance Function (SDF) [53, 54, 55, 56, 57] or Truncated Signed Distance Function (TSDF) [58, 59, 60, 61, 62, 63]. Nevertheless, the memory footprint can become excessive when high-resolution details are required.
Polygonal Mesh representations are composed of vertices and surfaces and describe complex 3D shapes compactly. However, their unstructured and non-differentiable nature [64] poses challenges in integrating them with neural networks for end-to-end differentiable pipelines. Some solutions addressing this problem, such as gradient approximation-based approaches [65, 66, 67, 68], can only use handcrafted gradient calculation. Other solutions like differentiable rasterizers [69, 70, 71] may lead to imprecise render results such as blurry content.
Neural Fields [72] have gained a growing interest in the 3D research community over recent years, departing from traditional representations reliant on geometric primitives [73]. A neural field is a mapping from spatial coordinates to scene properties (such as occupancy, color, radiance, etc), but unlike a voxel grid – in which the mapping is from a discrete cell to the value at that voxel – in a neural field the mapping is a learned function, typically a multi-layer perceptron. In this way neural fields implicitly learn compact, continuous, and differentiable 3D shape and scene representations.
One set of neural fields has focused on implicit surface representations. Occupancy Networks [74, 75, 76] encode shapes in a continuous 3D occupancy function represented by neural networks, using 3D point locations and features from point clouds, low-resolution voxels, or images to estimate occupancy probabilities. Concurrently, Deep SDF networks [77, 78, 79, 80] use neural networks to estimate SDFs from 3D coordinates and latent vectors. Recent methods like NeuS [79] and NeuS2 [80] notably improve surface reconstruction fidelity and efficiency for both static and dynamic objects.
Another set of approaches, referred to as Neural Radiance Fields (NeRF) [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94], display strong photorealistic rendering capabilities of the 3D world. These methods encode scene details with positional encoding techniques [95, 96, 90, 97], and utilize MLPs to predict radiance values (color and opacity) along camera rays. However, the necessity for MLPs to infer color and occupancy details for each sampled point in space (including those in empty spaces) requires significant computational resources. Thus, there is a strong incentive to reduce NeRFs’ computational overhead for real-time applications.
Hybrid Representations attempt to combine NeRF techniques with traditional volumetric-based approaches, facilitating high quality, real-time rendering [98, 99, 100, 101, 102, 103, 104, 105, 106, 107]. For instance, combining voxel grids [95, 96, 98, 97] or multi-resolution hash grids [102] with neural networks greatly reduces the NeRF’s training and inference time.
3D Gaussian Splatting [103, 106, 107] is a variation on point clouds where each point contains additional information representing the radiance being emitted in the spatial region around that point as an anisotropic 3D Gaussian ‘blob’. These 3D Gaussians are typically initialized from SfM point clouds [108], and optimized using differentiable rendering. 3D Gaussian Splatting achieves state-of-the-art novel-view synthesis at a fraction of NeRF’s computation by utilizing efficient rasterization [109] rather than raytracing.
2.2 Large Language Model (LLM)
@@
Traditional natural language processing (NLP) encompasses a wide range of tasks aimed at enabling systems to understand, generate and manipulate text. Early approaches in NLP relied on techniques like rule-based systems, statistical models, and earlier neural architectures like recurrent neural networks [110, 111]. The recent introduction of Large Language Models (LLMs), employing transformer architectures [95] and training on vast text corpora [112], has achieved unprecedented performance and sparked a new wave of excitement in the field. Since this paper focuses on 3D LLMs, we provide relevant background knowledge of LLMs here. For an in-depth exploration of LLMs at large, we refer to the latest surveys in the area [113, 114, 115, 116, 117].
@@ 2.2.1 LLM Architectures
In the context of LLMs, the “encoder-decoder” and “decoder-only” architectures are prominently used for NLP tasks.
Encoder-decoder architectures [95, 118, 119] consist of two main components: an encoder $f_{enc}$ and a decoder $f_{dec}$. The encoder and decoder components are typically implemented using transformers [95], which employ attention mechanisms to capture long-range dependencies in the input and output sequences. The encoder takes the input sequence $X=(x_{1},x_{2},\ldots,x_{N})$ and maps it into a sequence of latent representations $H=(h_{1},h_{2},\ldots,h_{N})$ that capture the contextual information, and the decoder generates the output sequence $Y=(y_{1},y_{2},\ldots,y_{T})$ based on $H$. Mathematically, the encoding process can be expressed as $H=f_{enc}(X)$, and the entire latent sequence $H$ is generated at once from $X$. The decoder, however, generates the output sequence $Y$ sequentially: $y_{t}=f_{dec}(y_{<t},H)$ where $y_{<t}=(y_{1},y_{2},\ldots,y_{t-1})$.
Decoder-only architectures [120, 121, 122], on the other hand, are a variant of the transformer architecture that uses only the decoder component. It is particularly suitable for language modeling tasks, where the goal is to predict the next token given the previous tokens. The decoder-only architecture can be mathematically expressed as $y_{t}=f_{dec}(y_{<t})$.
Tokenization is a preprocessing method to break the input text into a sequence of tokens, the basic data unit in language models. The number of tokens is finite, and each token can correspond to a word, sub-word, or a single letter. During inference, the input text is converted to a sequence of tokens and fed to the models, which predict the output tokens that are then converted back to text. The tokenization has a great impact on the performance of the language models, as it affects how the models perceive text. Various tokenization techniques are used, such as word-level tokenization, subword tokenization (e.g. byte-pair encoding [123], WordPiece [124], SentencePiece [125]), and character-level tokenization [126].
@@ 2.2.2 LLM Emergent Abilities
One major difference between LLM and traditional non-LLM methods is the emergent abilities that become available in large models but are not present in the small models [115]. The term “emergent abilities” refers to new, complex capabilities that arise as LLMs scale in size and complexity. These abilities enable advanced understanding and generation of natural language, problem-solving across various domains without specific training, and adapting to new tasks through in-context learning. In the following, we introduce several common emergent abilities in the scope of LLMs.
In-Context Learning  [127, 128, 129] refers to the ability of LLMs to understand and respond to new tasks or queries based on the context provided within the prompt, without requiring explicit retraining or fine-tuning. Milestone papers (GPT-2/GPT-3 [130, 127]) demonstrate in-context learning in a few-shot manner, where a model is provided with a few examples of a task within the prompt and then asked to tackle a different example without prior explicit training. State-of-the-art LLMs, such as GPT-4 [131], exhibit remarkable in-context learning abilities, understanding complex instructions, and performing a wide range of tasks from simple translations to generating code and creative writing, all based on the context provided within the prompt.
Reasoning in the context of LLMs, often referred to as “chain-of-thought” prompting [132, 133], involves the model generating intermediate steps or reasoning paths when tackling complex problems or questions. This approach allows LLMs to break tasks into smaller, manageable parts, facilitating a more structured and understandable solution process. To achieve this, training involves datasets that include a variety of problem-solving tasks [134, 135], logical puzzles [136, 137], and datasets designed to mimic reasoning under uncertainty [138]. Current state-of-the-art LLMs [127, 131, 139, 140, 141, 142, 143] demonstrate advanced reasoning capabilities typically when model sizes become larger than 60B to 100B parameters [132].
Instruction-following  [144, 127, 140, 145, 146] refers to the models’ capability to understand and execute commands, or follow instructions as specified by the user. This involves parsing the instruction, understanding its intent, and generating an appropriate response or action. Methods for adapting this ability to novel tasks may require instruction tuning from datasets [114] containing various instructions paired with correct responses or actions. Techniques such as supervised learning, reinforcement learning from human feedback, and interactive learning can further enhance performance.
@@ 2.2.3 LLM Fine-tuning
In the context of 3D-LLMs, LLMs are either utilized directly in their pre-trained state or undergo fine-tuning to accommodate novel multi-modal tasks. However, fine-tuning the entirety of an LLM’s parameters presents significant computational and memory challenges due to the large amount of parameters involved. Therefore, parameter-efficient fine-tuning (PEFT) [147] has become increasingly popular for adapting LLMs to specific tasks by updating only a relatively small subset of the model’s parameters instead of retraining the entire model. The following part lists four common PEFT methods used in LLMs.
Low-Rank Adaptation (LoRA) and variants [148, 149, 150] update parameters through low-rank matrices. Mathematically, a forward pass for LoRA [148] during the fine-tuning can be expressed as $h=W_{0}x+BAx$. $W_{0}$ is the frozen weight of LLM, while $BA$ is a low-rank matrix parameterized by newly introduced matrices $A$ and $B$, which are updated at fine-tuning stages. This method offers several distinct benefits. During fine-tuning, only $B$ and $A$ are optimized, significantly reducing the computational overhead associated with gradient calculation and parameter updates. Once the fine-tuning concludes and the weights are merged, there is no additional inference cost compared to the original model, as represented by the equation: $h=(W_{0}+BA)x$. Moreover, there’s no need to save multiple copies of LLMs for different tasks since multiple LoRA instances can be saved instead, reducing the storage footprint.
Layer Freezing [151, 142, 152] freezes selected layers of a pre-trained model while updating others during training. This often applies to layers closer to the model’s input or output, depending on the task’s nature and the model architecture. In 3D-LLM methods, for example, all layers except input and output embeddings [153] might be frozen to mitigate overfitting risks with task-specific datasets, retaining the pre-trained general knowledge and reducing parameters that need to be optimized.
Prompt Tuning [154, 155, 156] guides LLMs to perform specific tasks by framing them within prompts, adjusting model inputs in contrast to conventional fine-tuning which adjusts model parameters. Manual prompt engineering [132, 157, 158, 159] is the most intuitive method, yet it is difficult for experienced prompt-tuning engineers to find the optimal prompt. Another set of approaches is automatic prompt generation and optimization. One prevalent approach is to search for the exact optimal input prompt text, known as hard prompt, such as [160, 161, 162, 163]. Alternatively, one can use optimization approaches to optimize the embedding of the prompt (soft prompt[154, 164, 165, 166, 167, 168, 169, 170].
Adaptive Fine-tuning customizes the model architecture for specific tasks by adding or removing layers or modules [171, 172, 173]. This can include integrating new data modalities, such as visual information, alongside textual data. The core idea of adaptive fine-tuning is leveraging small neural network modules inserted between the layers of a pre-trained model. During adaptive fine-tuning, only the parameters of these adapter modules are updated, while the original model weights remain fixed.
2.3 2D Vision-Language models
Vision-Language Models are a family of models which aim to capture and exploit the relationship between text and images/video, and are capable of performing interactive tasks between the two modalities. The majority of VLMs have transformer-based architectures. By leveraging the attention module, visual and textual contents are conditioned on each other, which enables mutual interaction. In the following paragraphs, we briefly cover the applications of VLMs to discriminative and generative tasks.
Discriminative tasks involve predicting a certain characteristic about the data. VLMs, such as CLIP [174] and ALIGN [175], have shown exceptional performance in zero-shot transferability to unseen data in image classification. Both models comprise two modules: a visual encoder and a text encoder. Given an image and its category, CLIP and ALIGN are trained by maximizing the similarity between the image embedding and the textual embedding of the sentence “a photo of a {image category}”. The zero-shot transferability is achieved by replacing “{image category}” with possible candidates during inference and searching for sentences matching best with the image. These two works have inspired numerous follow-ups such as [176, 177, 178, 179], which further boost image classification accuracy. The models can also distill learned knowledge for other tasks, including object detection [180, 181, 182, 183], image segmentation [184, 185, 186, 187, 188], document understanding [189, 190] and video recognition [191].
Generative tasks utilize VLMs to generate text or images from input data. By leveraging large-scale training data, a single VLM can usually perform several image-to-text generation tasks, such as image captioning and visual question answering (VQA). Notable examples include SimVLM [192], BLIP [193], and OFA [194], etc. More powerful VLMs such as BLIP-2 [195], Flamingo [196], and LLaVA [197], are able to handle multi-turn conversations and reasoning based on the input image. Following the introduction of the diffusion model, text-to-image generation has also become a focus of the research community [198, 199]. By training on a large amount of image-text pairs, diffusion models [200, 201] can generate high-quality images based on textual input. This capability has also been extended to generating videos [202], 3D scenes [31], and dynamic 3D objects  [203]. Along with generative tasks, it is also possible to edit existing images through the textual prompt [204, 205, 206, 207].
2.4 Vision Foundation Models (VFMs)
Vision Foundation Models (VFMs) are large neural networks that aim to extract image representations that are sufficiently diverse and expressive to be directly deployed in various downstream tasks, mirroring the role pre-trained LLMs serve in downstream NLP tasks. One notable example is DINO [208], which uses a self-supervised teacher-student training paradigm. The representation learned achieves good results on both image classification and semantic image matching. The attention weights in DINO can also serve as segmentation masks for semantic components of the observed scene. Followup works like iBOT [209] and DINOv2 [210] further improve the representation by introducing the masked image modeling (MIM) loss. SAM, a transformer-based image segmentation model [211], is trained on a dataset consisting of 1.1 billion images with semantic masks and demonstrates strong zero-shot transfer capability. DINO (Zhang et al.[212] – not to be confused with DINO (Caron et al.[208] – employs a DETR-like [213] architecture and mixed query selection for object detection. The follow-up work Grounding-DINO [214] introduces text supervision to boost accuracy. Stable Diffusion [201], a text-to-image generator, has also been used as a feature extractor for ‘real’ images by running a single diffusion step on clean, or artificially noised images and extracting intermediate features [215, 169] or attention masks [216]. These features have been recently used for segmentation [216] and image matching [215, 169] tasks, due to the size and diversity of the training sets used for diffusion models, and due to observed emergent properties of diffusion features, such as zero-shot correspondence across images [215].

### v2

# When LLMs step into the 3D World: A Survey and Meta-Analysis of 3D Tasks via Multi-modal Large Language Models

@@ Abstract
As large language models (LLMs) evolve, their integration with 3D spatial data (3D-LLMs) has seen rapid progress, offering unprecedented capabilities for understanding and interacting with physical spaces. This survey provides a comprehensive overview of the methodologies that enable LLMs to process, understand, and generate 3D data. Highlighting the unique advantages of LLMs, such as in-context learning, step-by-step reasoning, open-vocabulary capabilities, and extensive world knowledge, we underscore their potential to significantly advance spatial comprehension and interaction within embodied Artificial Intelligence (AI) systems. Our investigation spans various 3D data representations, from point clouds to Neural Radiance Fields (NeRFs). It examines their integration with LLMs for tasks such as 3D scene understanding, captioning, question-answering, and dialogue, as well as LLM-based agents for spatial reasoning, planning, and navigation. The paper also includes a brief review of other methods that integrate 3D and language. The meta-analysis presented in this paper reveals significant progress yet underscores the necessity for novel approaches to harness the full potential of 3D-LLMs. Hence, with this paper, we aim to chart a course for future research that explores and expands the capabilities of 3D-LLMs in understanding and interacting with the complex 3D world. To support this survey, we have established a project page where papers related to our topic are organized and listed: https://github.com/ActiveVisionLab/Awesome-LLM-3D.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has marked a transformative era in natural language processing, enabling machines to understand, generate, and interact with human language in ways previously unimaginable. However, the physical world around us is inherently three-dimensional, and understanding spatial environments is crucial for many real-world applications that involve perception, navigation, and interaction within these 3D spaces. With recent advances, the application of LLMs has extended well beyond text. The fusion of LLMs with 3D data presents a unique opportunity to enhance computational models’ understanding of, and interaction with the physical world, leading to innovations across various domains, including autonomous systems [1, 2, 3, 4, 5], augmented reality [6, 7, 8, 9], robotic navigation [10, 11, 12], and robotic manipulation [13, 14, 15].
Recent works have demonstrated the potential of integrating LLMs with 3D data to interpret, reason, or plan in complex 3D environments, by leveraging the inherent strengths of LLMs, including zero-shot learning [16, 17], advanced reasoning [13, 18, 19], and extensive knowledge [20, 21]. However, despite rapid advances in this area, research remains scattered and relatively nascent, with most efforts emerging only since 2023 (see Fig. 1). This fragmentation creates a need for a comprehensive and systematic survey to consolidate the current state of research and provide a clear overview of the methodologies, challenges, and future directions. The aim of this survey is to fill these gaps by providing a detailed analysis of the field, enabling researchers to better understand the progress made thus far and to identify key areas for future exploration.
To the best of our knowledge, we are the first to survey the field of 3D-LLMs. We outline the methodologies that are currently being used to allow LLMs to ingest and process 3D information in Fig. 3, and present a novel taxonomy categorizing 3D-LLM methods based on the role of the language model in solving 3D tasks in Fig. 4. We describe areas where the amalgamation of LLMs and 3D data has been successfully demonstrated, such as: using LLMs’ world-knowledge [20, 22] and reasoning capabilities [21, 23] to enhance 3D task performance, using LLMs as multi-modal interfaces [24, 18] and embodied agents [15, 13], or generating complex 3D scenes with LLMs [25, 26]. We then provide a multifaceted comparison between existing methods, with a focus on the 3D vision components, LLM components, and alignment methodologies used to train these models (see Tab. I).
Figure 1: An overall timeline to show the developments of 3D representations (Sec. 2.1), L(V)LMs (Sec. 2.2 and Sec. 2.3), 3D w/VLMs&VFMs(Sec. 5), 3D w/LLMs (Sec. 4). This figure illustrates how the development of 3D representation and L(V)LMs inspired the 3D vision-language methods (3D w/LLMs, 3D w/VLMs VFMs). Besides, we can clearly see the rapid growth of 3D-LLM methods starting from 2023, which calls for the attention of more researchers.
Compared to other LLM-related surveys, which either focus on single tasks like video understanding [27]—or on specific scenarios like robotics [28, 29, 30] or autonomous driving [31], our survey focuses on the broad potential of integrating LLMs with 3D data. Based on this, we highlight how the advantages and disadvantages of different 3D representations (Sec. 2.1), and how these representations can be combined with LLMs to accomplish various 3D tasks (Sec. 4). Additionally, we cover 3D methods utilizing Vision-Language models (VLMs) and Vision-Foundation Models (VFMs) (Sec. 5). Altogether, these solutions could provide insights for various real-world scenarios such as AR/VR, healthcare, game development, and interior design.
In addition to the above, we offer a thorough review of 3D vision-language tasks (Sec. 3), categorizing them into five major types based on differences in the input and output modalities. We also compare datasets used for training and testing (Sec. 6), mapping the trajectory of this field through the evolution of datasets (see Fig. 5). Finally, we identify key challenges and outline future directions (Sec. 7).
@@ 2 Background
This section offers background knowledge on 3D representations, Large Language Models (LLMs), 2D Vision-Language Models (VLMs), and Vision Foundation Models (VFMs).
@@
Choosing an effective 3D representation is a critical design decision in 3D-LLMs research. It is also a fundamental research area in computer vision and graphics, driven by advances in deep learning, computational resources, and the availability of 3D data. In practice, image-based representations, such as multi-view RGB images and RGB-D data, although not strictly 3D, are often employed as proxies for 3D geometry, owing to their ease of acquisition and strong compatibility with existing 2D vision backbones. Nevertheless, these representations mainly capture geometry at the pixel level. Conversely, structured scene representations, such as 3D scene graphs, offer semantic abstraction over spatial layouts and are increasingly adopted for reasoning tasks in embodied AI. Below, we briefly describe the most common 3D representations.
Point Clouds represent 3D shapes as discrete data points in space, typically storing the XYZ coordinates and optionally additional attributes such as color or surface normals. Point cloud-based methods [32, 33] are lightweight and easy to obtain (e.g., from LiDAR, structured light, or photogrammetry), but lack explicit surface topology.
Voxel Grids [34, 35] divide the 3D space into regular volumetric units (voxels), analogous to pixels in 2D. Each voxel can encode occupancy (binary or probabilistic), surface distance (e.g. Signed Distance Function [36, 37]), or Truncated SDF [38, 39]. While intuitive, voxel grids become memory-intensive at high resolution, which motivates sparse voxel structures [40].
Meshes use interconnected vertices, edges, and faces to compactly model 3D geometry with surface continuity. Despite their expressiveness, they are less favored in learning-based pipelines due to their irregular topology. Solutions include gradient approximation [41, 42] and differentiable rasterization [43], though these often suffer from rendering artifacts or limited precision.
Neural Fields [44] represent 3D scenes as continuous functions (typically MLPs) that map spatial coordinates to physical properties such as occupancy, color, or semantics. Unlike discrete voxel grids, neural fields model scenes as continuous volumes. For example, Occupancy Networks [45, 46] predict whether a point lies inside an object, while Deep SDF [47, 48] learns continuous signed distance fields. These representations are compact and flexible but often slow to evaluate.
Neural Radiance Fields (NeRF) [49, 50] extend neural fields by modeling view-dependent color and density along camera rays. This enables photorealistic novel-view synthesis but requires expensive sampling and evaluation, especially in empty space, prompting research on acceleration.
Hybrid Representations combine neural fields with explicit structures to improve efficiency. For instance, voxel grids [51] or multi-resolution hash grids [52] are used to condition lightweight MLPs, significantly reducing NeRF training and inference cost while maintaining high rendering quality.
3D Gaussian Splatting [53] offers a differentiable point-based representation, where each point emits anisotropic Gaussian blobs instead of sharp vertices. This enables real-time, high-fidelity rendering. Initializations typically come from SfM-based point clouds [54], though recent works [55, 56] explore SfM-free pipelines.
Figure 2: Common task categories for 3D Vision Language models. Each category contains many types of queries, but we provide a single example from each category. The scan is from the ScanNet dataset [57], and the ‘question answering’ and ‘situated question answering’ examples are adapted from ScanQA [58] and SQA3D [59] respectively.
2.2 Large Language Models (LLMs)
@@
Natural Language Processing (NLP) aims to enable systems to understand, generate, and manipulate text. Early approaches used rule-based systems, statistical models, or neural architectures like recurrent neural nets [60]. The rise of LLMs, powered by transformers [61] and vast text corpora [62], has transformed the field. Since this paper focuses on 3D LLMs, we provide a brief background on LLMs here. For a more detailed coverage, see recent surveys on this topic [63, 64, 65].
@@ 2.2.1 LLM Architectures
Large language models are typically built upon either encoder-decoder or decoder-only transformer architectures [61]. Encoder-decoder models process the input sequence via an encoder to produce contextual embeddings, which are then consumed by the decoder to autoregressively generate the output. In contrast, decoder-only architectures [66] generate outputs directly from previous tokens without a separate encoder, making them particularly efficient for generative tasks. Another critical component is tokenization, which converts raw text into discrete tokens. Choices like byte-pair encoding (BPE) [67], WordPiece [68], and SentencePiece [69] significantly impact model performance and vocabulary expressivity. These architectural and preprocessing designs influence how language and multimodal inputs are represented and processed, and they are often adapted or extended in 3D-LLMs to support spatial encoders and cross-modal alignment.
@@ 2.2.2 LLM Emergent Abilities
A defining feature of LLMs is their emergent abilities, which arise as model scale increases [65], enabling zero-shot generalization across diverse tasks without explicit training. These include in-context learning [70], where the model adapts to new tasks based on prompts alone; reasoning, often enabled by chain-of-thought prompting [71], where intermediate steps are generated to solve complex problems; and instruction-following [72], where the model executes natural language commands via instruction tuning. These abilities underpin the remarkable generality of LLMs and have directly inspired recent 3D-LLMs to adopt similar paradigms—using in-context demonstrations, compositional reasoning, and open-ended instruction-following over multimodal 3D inputs to enhance task generalization without task-specific retraining.
@@ 2.2.3 LLM Fine-tuning
Given the large number of parameters in LLMs, full fine-tuning is often infeasible for 3D-LLMs. Instead, parameter-efficient fine-tuning (PEFT) methods have become the standard to adapt LLMs to 3D tasks with reduced computational cost. LoRA [73] and its variants [74, 75] insert trainable low-rank matrices into transformer layers, allowing efficient adaptation while keeping most of the original weights frozen. Layer freezing [76, 77] updates only a small subset of layers, typically at the input/output, preserving general language knowledge and reducing overfitting. Prompt tuning steers model behavior by modifying input prompts instead of model parameters, including handcrafted prompts [71], discrete token search (hard prompts) [78], or learned embeddings (soft prompts) [79]. Finally, adaptive tuning [80, 81] integrates lightweight neural modules between frozen layers to enable efficient cross-modal adaptation, such as incorporating 3D data alongside text. These PEFT strategies are widely adopted in 3D-LLMs to enable scalable and modular alignment of LLMs with 3D inputs.
2.3 Vision-Language and Vision Foundation Models
2D Vision and Vision-Language Foundation Models aim to learn universal visual representations and multimodal alignments, enabling transfer to a wide range of downstream tasks.
Vision-Language Models (VLMs) such as CLIP [82] and ALIGN [83] align image and text embeddings through contrastive training, achieving strong zero-shot transfer in classification and extending to tasks like detection [84], segmentation [85], and video understanding [86]. Later VLMs like BLIP-2 [87], Flamingo [88], and LLaVA [89] support image-conditioned generation and multi-turn interaction. Text-to-image diffusion models [90] further expand the generative capabilities, with extensions to video [91] and 3D generation [92, 93].
Vision Foundation Models (VFMs) focus on learning rich visual representations via self-supervised or large-scale supervised pretraining. DINO [94] and its successors [95, 96] use contrastive and masked image modeling strategies, demonstrating generalization in classification, segmentation, and matching. Segment Anything Model (SAM)[97] enables zero-shot segmentation across diverse domains. Beyond this, diffusion models[90] also serve as feature extractors [98, 99], revealing semantic priors through intermediate representations.
These 2D VLMs and VFMs lay the foundation for recent efforts extending multimodal learning into the 3D domain, which we discuss in the following sections.